OpenStudy (anonymous):

Need Some assistance Paul predicts that n stalks of corn planted together will produce (2n squared - 2n) ears of corn. If paul plants 3 corn plants, how many ears of corn does paul expect to harvest?

OpenStudy (anonymous):

\[2n^2-2n=2n(n-1)\] and he plants n = 3 get \[2\times 3(3-1)=6\times 2=12\] ears

OpenStudy (anonymous):

or if you prefer \[2\times 3^2-2\times 3=18-6=12\]

OpenStudy (anonymous):

ok thank you so much